OpenStudy (anonymous):

y=-5 and x=3 Write a variation equation for both.

OpenStudy (anonymous):

Y varies directly as x means y = kx. So, plug in y = -18 and x = -2 to find k -18 = k(-2) -18 = -2k 9 = k So, your general model that will work for any x and y is: y = 9x. If x is 20, then y = 9(20) = 180. For your first question. y varies inversly as x means y = k/x So, plug these in again to find k. 5 = k/3 15 = k. y = 15/x

OpenStudy (anonymous):

There also is another problem that goes along with it. x:1 5 10 and 15 y: 30 6 3 and 2.

OpenStudy (anonymous):

i dont under stand

OpenStudy (anonymous):

Uhm, okay for the second question it wants you circle the number that was the constant? I don't know im confused myself..